# A bottleneck model with shared autonomous vehicles: Scale economies and price regulations

### v1

@@ 5.4 Regulatory strategy lowering social cost while achieving a Pareto improvement
Hence, the regulatory strategy proposed in Section 3 can improve both the equilibrium commuting cost and the social cost.
@@
Promoting SAV usage through AC pricing still reduces commuting costs, yet it increases the social cost. The reason is that a small $\eta$, reflecting a weak capacity effect $\kappa$ with a strong VOT effect $\theta$, provides only limited social gains from further SAV adoption. For large $N$, a relatively high AC fare encourages additional SAV adoption; however, the congestion created among SAV commuters outweighs the benefit to NV commuters. Although the fare reduction offsets this negative effect, it still induces a decrease in the service provider’s profit; as a result, the social cost increases even though commuting costs decreases.
The discussion suggests that switching price regulation solely once the share of SAV commuters exceeds a threshold can raise the social cost. If the concern is confined to commuters’ costs, introducing the price regulation according to this single criterion is adequate because it surely reduces the commuter’s costs. By contrast, if a policymaker would like to additionally reduce social cost (i.e., increase the providers’ profit), the threshold rule can be supplemented with a second criterion: the technological maturity of the capacity effect, represented by an increase in $\eta$. This combination offers an option that can simultaneously reduce both commuting costs and the social cost. Specifically, we propose the following regulatory strategy:
@@
No price regulation. Keep fares unregulated and allow a natural monopoly.
@@
Technological maturation under monopoly. Maintain the natural monopoly until the index satisfies $\eta\geq 1$. Operationally, this means waiting until the capacity-expansion parameter $\kappa$ has fallen sufficiently.
@@
Activate AC pricing. Once $\eta\geq 1$ is satisfied, impose AC pricing.
@@
This pricing strategy suggests a simple rule: postpone price regulation until the capacity-expansion technology has matured, for example, until SAV platooning can reliably sustain shorter headways. This rule eliminates the discrepancy between reductions in equilibrium commuting costs and social costs. In the early diffusion stage, the capacity effect would be weak ($\eta$ is low), so allowing a monopoly is socially preferable. During this period, resources should be channelled toward accelerating technological advances that strengthen capacity expansion. Once the technology matures and SAV adoption expands, switching to AC pricing becomes socially beneficial for SAV commuters; both the equilibrium cost and the social cost decrease.
@@
Figure 6 shows the same social-cost contour plot as in Figure 5, now used to illustrate the effectiveness of the proposed pricing strategy. Assume first that $\eta=0.15$. If a natural monopoly is allowed, the number of SAV commuters rises to the monopoly‐equilibrium level discussed in Section 3 (green circle in the figure). Introducing AC pricing at that point would raise the SAV share; as indicated by the blue circle with a broken outline, the social cost would increase even though the SAV share increases. Next, follow the strategy in Phase 2 and continue with the natural monopoly until $\eta$ becomes $1.2$. The social-cost contours indicate that the traffic state has now shifted into a region where the capacity-expansion effect is strong enough for additional SAV commuters to reduce social cost. When the system reaches this state, switching to AC pricing (Phase 3) results in a further increase in the number of SAV commuters, while simultaneously reducing the social cost (shifting to the blue circle).
In summary, our results identify $\eta$ as the key parameter that aligns reductions in social cost with reductions in equilibrium cost in both first-best and second-best policy design. This finding underscores the importance of allocating limited innovation budgets to capacity-expansion technologies that deliver social benefits. Therefore, effective utilization of SAVs in the transport system requires price regulation that is tied not only to the current level of SAV adoption but also to the maturity of the underlying technology; combining these two triggers is essential for realizing the full social benefit of SAVs.
@@ 6 Conclusion
This paper examines how scale economies in SAV operations affect the efficiency of a mixed-traffic transportation system by analyzing three fare-setting scenarios: MC pricing, AC pricing, and a natural-monopoly outcome. First, our equilibrium analysis showed that MC pricing, although it forces the operator to run at a deficit, yields the lowest commuting cost among the three scenarios. We also demonstrated a two-step regulatory strategy: allow a natural monopoly initially and switch to AC pricing once SAV adoption has passed a critical threshold. This strategy eliminates the deficit while further promoting SAV use and reducing commuting costs. In addition, we find that the Downs–Thomson paradox can arise in SAV system under AC pricing: capacity expansion can raise commuting costs because the SAV share falls. These findings confirm the superiority of MC pricing for reducing commuting costs, while also demonstrating that a well-timed implementation of AC pricing can achieve a convenient and economically sustainable transportation system for commuters.
@@
Next, we evaluate social cost, defined as the sum of commuters’ total costs and the operator’s profit. When a time-varying congestion toll can be implemented alongside fare regulation, the first-best policy is MC pricing combined with the optimal toll. Using the closed-form solution, we derive the condition under which implementing the optimal toll to the MC equilibrium produces a Pareto improvement: the SAV capacity effect must exceed its VOT effect. We also propose that the toll revenue is used for financing the road capacity and show that the self-financing principle holds. We then solve the second-best problem, in which only the fare can be regulated, and compare its solution with the three benchmark scenarios. The results show that MC pricing does not always minimize the social cost. When the capacity effect is weaker than the VOT effect, promoting SAV use can increase congestion externalities; their impact is reflected in a decline in the operator’s profit. In the extreme case of a negligible capacity effect, the natural-monopoly outcome becomes the most efficient. Overall, the impact of fare regulation can become far from intuitive. The choice and timing of MC or AC pricing must be tailored to the current SAV adoption level and the maturity of the underlying technology to achieve a socially efficient transportation system.
@@
A natural direction for future research is to extend our model spatially, for example, to a corridor or network with multiple bottlenecks. In such a setting, it is important to examine the spatio-temporal patterns of NV and SAV use and to assess how fare regulation alters those patterns. It would also be important to examine the flow and cost patterns when commuters can choose their residential locations in the long run. The literature shows that commuters sort themselves spatially according to their heterogeneity (e.g., Takayama and Kuwahara, 2017; Osawa et al., 2018; Sakai et al., 2024). Hence, adding a mode-choice dimension is expected to yield a clear sorting pattern in which each mode is preferred by commuters living at specific distances from the common destination. That pattern could then be used to derive policy insights not only for fare regulation but also for land-use controls. egarding mode choices, it is important to incorporate public transit. Because transit operations would be subject to different scale economies, their inclusion is likely to generate additional multiple equilibria with varying shares of NV, SAV, and transit commuters. Analyzing them would clarify how best to combine SAV services with conventional transit. It would also be interesting to explore new schemes for redistributing the service provider’s profit. This study shows that the second-best policy can coincide with profit maximization by a monopolistic provider; accordingly, an appropriate redistribution of those profits is crucial for broad social acceptance of the policy. Finally, it would be worthwhile to explore alternative market structures, such as oligopolistic competition among SAV operators. These analyses deepen our understanding of SAV systems that incorporate scale economies.
@@ Acknowledgement
This work was supported by JSPS KAKENHI Grant Numbers JP23K13418 and JP25K01339.
@@ Appendix A Closed-form solution of mode-choice equilibrium
$\cfrac{\beta\gamma}{\mu(\beta+\gamma)}(N_{n}+\kappa N_{a})+t_{f}+F_{n}=\cfrac{\beta\gamma}{\mu(\beta+\gamma)}(\theta N_{n}+\kappa N_{a})+\theta t_{f}+p.$ (87)
@@
$\cfrac{\beta\gamma(1-\theta)}{\mu(\beta+\gamma)}(N-N_{a})-(p-m)=\theta t_{f}-t_{f}+m-F_{n}.$ (88)
@@ A.1 MC pricing
$N_{a}^{MC*}=N-\cfrac{\mu(\beta+\gamma)}{\beta\gamma(1-\theta)}(\theta t_{f}-t_{f}+m-F_{n})=\cfrac{AN-B}{A}.$ (89)
@@ A.2 AC pricing
Consider the case where $N_{a}\neq 0$. Substituting the equation into Eq. (88) yields the quadratic equation,
@@ Appendix C Temporal-sorting property under the first-best policy
$s(\bar{t})+t_{f}+F_{n}+\tau(\bar{t})=c^{FB*},\qquad s(\bar{t})+\theta t_{f}+m+\kappa\tau(\bar{t})=c^{FB*}.$
@@
$(1-\kappa)\,s(\bar{t})=(1-\kappa)c^{FB*}-\bigl{[}\theta t_{f}+m-\kappa(t_{f}+F_{n})\bigr{]},$
so $s(\bar{t})$ must equal a constant. Because $s(\cdot)$ is strictly convex and continuous, the set of times at which it attains the same value has measure zero. Hence $\mathcal{T}$ cannot contain an interval of positive measure.
@@
$\left[s(t)+\theta t_{f}+m+\kappa\tau(t)\right]-\left[s(t)+t_{f}+F_{n}+\tau(t)\right]=\begin{cases}>0,\quad&\text{if}\quad\tau(t)<\cfrac{B}{1-\kappa},\\ 0,\quad&\text{if}\quad\tau(t)=\cfrac{B}{1-\kappa},\\ <0,\quad&\text{if}\quad\tau(t)>\cfrac{B}{1-\kappa}.\end{cases}$ (122)
@@
    $\left[s(t)+\theta t_{f}+m+\kappa\tau(t)\right]-\left[s(t)+t_{f}+F_{n}+\tau(t)\right]\geq 0.$ (123)
@@
    $\left[s(t)+t_{f}+F_{n}+\tau(t)\right]-\left[s(t)+\theta t_{f}+m+\kappa\tau(t)\right]>0.$ (125)
    We thus have $\tau(t)>B/(1-\kappa)$. We also have $s(t)+\theta t_{f}+m+\kappa\tau(t)=c^{FB*}$. Therefore,
@@ D.1 Case (i): mixture of normal and shared autonomous vehicles
where $t_{n}^{-}$ and $t_{n}^{+}$ are the earliest and latest destination-arrival times of NV commuters; $t_{a}^{-}$ and $t_{a}^{+}$ are those of SAV commuters.
@@
The patterns of destination-arrival flow of NV and SAV commuters are given as follows:
@@
$s(t)+\theta t_{f}+m+\kappa\tau(t)>s(t)+t_{f}+F_{n}+\kappa\tau(t)\geq c^{FB*}+\kappa\tau(t)\geq c^{FB*}.$ (133)
@@
$s(t)+\theta t_{f}+m+\kappa\tau(t)=s(t)+\theta t_{f}+m<c^{FB*}+B-\cfrac{B}{1-\kappa}<c^{FB*}.$ (138)
@@
Using the destination-arrival flow and the earliest/latest destination-arrival times of NV and SAV commuters, their total numbers $N_{n}$ and $N_{a}$ are written as follows:
@@
Finally, substituting $c^{FB*}$ into the equation, we have the earliest/latest destination-arrival times, as follows:
@@ D.2 Case (ii): normal vehicles only
where $t^{-}$ and $t^{+}$ denote the earliest and latest destination-arrival times of commuters, respectively. They are represented as
@@ F.1.2 Comparison between $SC^{MC*}$ and $SC_{0}^{AC*}$
F.2 MC pricing vs. natural monopoly
@@
F.3 AC pricing vs. natural monopoly

### v2

@@ 5.4 Pricing strategy for social-cost reduction and Pareto improvement
Hence, the pricing strategy can improve both cost measures.
@@
Promoting SAV adoption through AC pricing still reduces commuting costs, yet it increases the social cost. The reason is that a small $\eta$, reflecting a weak capacity effect $\kappa$ with a strong VOT effect $\theta$, provides only limited social gains from further SAV adoption. Switching to AC pricing increases the number of SAV commuters (see Section 3.3). However, because the capacity effect is weak, this increase worsens congestion among SAV commuters, as discussed in Section 5.2. While the negative impact on commuters’ costs is offset by the fare reduction under AC pricing, this fare reduction, in turn, lowers the service provider’s profit. As a result, the social cost can increase even when commuting costs decrease.
The discussion suggests that a pricing strategy that switches between fare-setting rules based solely on SAV adoption can increase the social cost. If a policymaker’s concern is confined to commuters’ costs, such a strategy is sufficient, as it reduces them. By contrast, if the policymaker also aims to reduce the social cost, this threshold rule can be augmented with a second criterion: the technological maturity of the capacity effect, captured by $\eta$. Specifically, we propose the following pricing strategy:
@@
No fare regulation. Keep the fare unregulated and allow monopoly pricing.
@@
Technological maturation under monopoly. Maintain monopoly pricing until $\eta\geq 1$ holds. Operationally, wait until $\kappa$ has fallen sufficiently.
@@
Introduce AC fare regulation. Once $\eta\geq 1$ is satisfied, implement AC pricing.
@@
This pricing strategy suggests a simple rule: postpone fare regulation until the capacity-expansion technology has matured, for example, until SAV platooning can reliably sustain shorter headways. In the early diffusion stage, the capacity effect would be weak ($\eta$ is low), so allowing unregulated monopoly pricing is socially preferable. During this period, resources should be channelled toward accelerating technological advances that strengthen capacity expansion. Once the technology matures and SAV adoption expands, switching to AC pricing becomes socially beneficial for SAV commuters.
@@
Figure 6 shows the same social-cost contour plot as in Figure 5, now used to illustrate the effectiveness of the proposed pricing strategy. Assume first that $\eta=0.15$. Under unregulated monopoly pricing, the number of SAV commuters increases to the monopoly-equilibrium level (the green circle in the figure). Switching to AC pricing increases the number of SAV commuters further (the blue circle with a dashed outline); however, it raises the social cost. Next, we follow Phase 2 of the proposed strategy and maintain unregulated monopoly pricing until $\eta$ becomes $1.2$. The figure indicates that the traffic state has shifted into a region where the capacity-expansion effect is strong enough that additional SAV commuters reduce the social cost. At this point, switching to AC pricing (Phase 3) further increases the number of SAV commuters while simultaneously reducing the social cost (shifting to the blue circle).
In summary, our results identify $\eta$ as the key parameter that aligns reductions in social cost with reductions in equilibrium cost in both first-best and second-best policy design. This finding underscores the importance of allocating limited innovation budgets to capacity-expansion technologies that deliver social benefits. Therefore, effective utilization of SAVs in the transport system requires fare regulation that is tied not only to the current level of SAV adoption but also to the maturity of the underlying technology; combining these two triggers is essential for realizing the full social benefit of SAVs.
@@ 6 Conclusion
This paper examines how scale economies in SAV operations affect the efficiency of a mixed-traffic transportation system by analyzing three fare-setting scenarios: MC pricing, AC pricing, and unregulated monopoly pricing. First, our equilibrium analysis showed that MC pricing, although it forces the operator to run at a deficit, yields the lowest commuting cost among the three scenarios. We also demonstrated a two-step pricing strategy: allow monopoly pricing initially and switch to AC pricing once SAV adoption has passed a critical threshold. This strategy eliminates the deficit while promoting SAV adoption and reducing commuting costs. In addition, we find that the Downs–Thomson paradox can arise in SAV system under AC pricing: capacity expansion can increase commuting costs because the SAV adoption decreases. These findings confirm the superiority of MC pricing for reducing commuting costs, and demonstrate that a well-timed implementation of AC pricing can achieve a convenient and economically sustainable transportation system for commuters.
@@
Next, we evaluate the social cost, defined as the commuters’ total costs minus the operator’s profit. When a time-varying congestion toll can be implemented alongside fare regulation, the first-best policy is MC pricing combined with the optimal toll. Using the closed-form solution, we derive the condition under which implementing the optimal toll to the MC equilibrium produces a Pareto improvement: the SAV capacity effect must exceed its VOT effect. We also propose that the toll revenue is used for financing the road capacity and show that the self-financing principle holds. We then solve the second-best problem, in which only the fare can be regulated, and compare its solution with the three fare-setting scenarios. The results show that MC pricing does not always minimize the social cost. When the capacity effect is weaker than the VOT effect, promoting SAV adoption can increase congestion externalities; their impact is reflected in a decrease in the operator’s profit. In the extreme case of a negligible capacity effect, unregulated monopoly pricing becomes the most efficient. Overall, the impact of fare regulation can become far from intuitive. The choice and timing of MC or AC pricing must be tailored to the current SAV adoption level and the maturity of the underlying technology to achieve a socially efficient transportation system.
@@
A natural direction for future research is to extend our model spatially, for example, to a corridor or network with multiple bottlenecks. In such a setting, it is important to examine the spatio-temporal patterns of NV and SAV use and to assess how fare regulation alters those patterns. It would also be important to examine flow and cost patterns in long-run settings where travel demand may be elastic, for example, when commuters can adjust their residential locations and when the number of commuters is endogenously determined. The literature shows that commuters sort themselves spatially according to their heterogeneity (e.g., Takayama and Kuwahara, 2017; Osawa et al., 2018; Sakai et al., 2024). Hence, adding a mode-choice dimension is expected to yield a clear sorting pattern in which each mode is preferred by commuters living at specific distances from the common destination. In addition, allowing total demand $N$ to respond to generalized travel cost (e.g., Arnott et al., 1993; Fosgerau et al., 2018) would be important for capturing potential induced-demand effects when SAVs lower travel costs. These extensions could then provide policy insights not only for fare regulation but also for land-use policy. Incorporating an explicit parking model that accounts for parking location choice and capacity constraints (Arnott and Inci, 2010; Liu, 2018; Tian et al., 2019a) is another promising direction. This would link NV parking costs to traffic conditions and provide a more refined assessment of NV–SAV adoption and the welfare impacts of fare regulation.
On the supply side, an important direction for future research is to explicitly incorporate SAV operational features, for example, by modeling pickup waiting time and deadheading as outcomes of an SAV operator’s fleet decisions and operations. Such operational frictions can worsen congestion and, in some cases, may reduce effective bottleneck capacity as SAV usage increases (i.e., $\kappa>1$). Because these supply-side mechanisms may interact with the scale economies considered in our framework, it would be important to examine whether, and to what extent, the equilibrium outcomes and policy implications identified here are preserved under such extensions. Explicitly modeling the operator’s fleet decisions is also closely related to a key advantage of SAVs: vehicles can be redeployed for other tasks after drop-off. It would therefore be interesting to incorporate this feature explicitly and to assess the impacts of SAV adoption from this perspective. Another important direction is to incorporate public transit. Because transit operations would be subject to different scale economies, their inclusion is likely to generate additional multiple equilibria with varying shares of NV, SAV, and transit commuters. Analyzing them would clarify how best to combine SAV services with conventional transit. It would also be interesting to explore new schemes for redistributing the service provider’s profit. This study shows that the second-best policy can coincide with profit maximization by a monopolistic provider; accordingly, an appropriate redistribution of those profits is crucial for broad social acceptance of the policy. Finally, it would be worthwhile to explore alternative market structures, such as oligopolistic competition among SAV operators. These analyses deepen our understanding of SAV systems that incorporate scale economies.
@@ Acknowledgement
We thank Minoru Osawa for their helpful comments. We are also grateful to the associate editor and the anonymous referees for their valuable comments. This work was supported by JSPS KAKENHI (Grants JP23K13418, JP25K01339, and JP25H00751), JST FOREST Program (Grant JPMJFR215M), and the Council for Science, Technology and Innovation (Grant JPJ012187).
@@ Appendix A Closed-form solution of mode-choice equilibrium
$\cfrac{\beta\gamma}{\mu(\beta+\gamma)}(N_{n}+\kappa N_{a})+t_{f}+F_{n}=\cfrac{\beta\gamma}{\mu(\beta+\gamma)}(\theta N_{n}+\kappa N_{a})+\theta t_{f}+p+w.$ (89)
@@
$\cfrac{\beta\gamma(1-\theta)}{\mu(\beta+\gamma)}(N-N_{a})-(p-m)=\theta t_{f}+m+w-t_{f}-F_{n}.$ (90)
@@ A.1 MC pricing
$N_{a}^{MC*}=N-\cfrac{\mu(\beta+\gamma)}{\beta\gamma(1-\theta)}(\theta t_{f}+m+w-t_{f}-F_{n})=\cfrac{AN-B}{A}.$ (91)
@@ A.2 AC pricing
Consider the case where $N_{a}>0$. Substituting the equation into Eq. (90) yields the quadratic equation,
@@ Appendix C Temporal-sorting property under the first-best policy
$s(\bar{t})+t_{f}+F_{n}+\tau(\bar{t})=c^{FB*},\qquad s(\bar{t})+\theta t_{f}+m+w+\kappa\tau(\bar{t})=c^{FB*}.$
@@
$(1-\kappa)\,s(\bar{t})=(1-\kappa)c^{FB*}-\bigl[\theta t_{f}+m+w-\kappa(t_{f}+F_{n})\bigr],$
so $s(\bar{t})$ must equal a constant. Because $s(\cdot)$ is continuous and strictly monotone on each of $(-\infty,0]$ and $0,\infty)$, for any $\ell\in\mathbb{R}$ the level set $\{t\in\mathbb{R}\mid s(t)=\ell\}$ contains at most one point in each half-line, and hence at most two points in total. Therefore, it has Lebesgue measure zero, and $\mathcal{T}$ cannot contain an interval of positive measure.
@@
$\left[s(t)+\theta t_{f}+m+w+\kappa\tau(t)\right]-\left[s(t)+t_{f}+F_{n}+\tau(t)\right]=\begin{cases}>0,\quad&\text{if}\quad\tau(t)<\cfrac{B}{1-\kappa},\\ 0,\quad&\text{if}\quad\tau(t)=\cfrac{B}{1-\kappa},\\ <0,\quad&\text{if}\quad\tau(t)>\cfrac{B}{1-\kappa}.\end{cases}$ (124)
@@
    $\left[s(t)+\theta t_{f}+m+w+\kappa\tau(t)\right]-\left[s(t)+t_{f}+F_{n}+\tau(t)\right]\geq 0.$ (125)
@@
    $\left[s(t)+t_{f}+F_{n}+\tau(t)\right]-\left[s(t)+\theta t_{f}+m+w+\kappa\tau(t)\right]>0.$ (127)
    We thus have $\tau(t)>B/(1-\kappa)$. We also have $s(t)+\theta t_{f}+m+w+\kappa\tau(t)=c^{FB*}$. Therefore,
@@ D.1 Case (i): mixture of normal and shared autonomous vehicles
where $t_{n}^{-}$ and $t_{n}^{+}$ are the earliest and latest destination arrival times of NV commuters; $t_{a}^{-}$ and $t_{a}^{+}$ are those of SAV commuters.
@@
The patterns of destination arrival flow of NV and SAV commuters are given as follows:
@@
$s(t)+\theta t_{f}+m+w+\kappa\tau(t)>s(t)+t_{f}+F_{n}+\kappa\tau(t)\geq c^{FB*}+\kappa\tau(t)\geq c^{FB*}.$ (135)
@@
$s(t)+\theta t_{f}+m+w+\kappa\tau(t)=s(t)+\theta t_{f}+m+w<c^{FB*}+B-\cfrac{B}{1-\kappa}<c^{FB*}.$ (140)
@@
Using the destination arrival flow and the earliest/latest destination arrival times of NV and SAV commuters, their total numbers $N_{n}$ and $N_{a}$ are written as follows:
@@
Finally, substituting $c^{FB*}$ into the equation, we have the earliest/latest destination arrival times, as follows:
@@ D.2 Case (ii): normal vehicles only
where $t^{-}$ and $t^{+}$ denote the earliest and latest destination arrival times of commuters, respectively. They are represented as
@@ F.1.2 Comparison between $SC^{MC*}$ and $SC_{0}^{AC*}$
F.2 MC pricing vs. unregulated monopoly pricing
@@
F.3 AC pricing vs. unregulated monopoly pricing